Hey, I really need help it this question:
A car leaves a country town at 60km per hour. Two hours later, a second car leaves town; it catches the first car after 5 more hours. Find the speed of the second car.

Answer by ewatrrr(24785)


Hi,
*Note: both cars will have traveled the same distance. d = r*t
second car takes 7hrs. first car takes 5hrs to travel this distance
letting x reperesent the speed of the second car
60mph*7hr = x * 5hr
60mph*7/5=x
x = 84mph